# Human Attention During Localization of Memory Bugs in C Programs

### v1

@@ 3.6 Data Processing
We graded the participants’ answers on a scale of 1-5 where 1 was very inaccurate and 5 was highly accurate. Attempts without answers received 0. A score of highly-accurate meant that the reported location was identical to the one in the official patch fix or we determined to be another fix of equal or higher quality. To minimize bias, we applied the rubrics shown in Tables II and III to decide scores.
@@
TABLE II: Rubric for grading “what” questions.
@@
TABLE III: Rubric for grading “where” questions.
@@
To process and analyze the data, we used the “Map Tokens” and “Generate Fixations” tools in iTrace Toolkit V0.2.2 to extract fixations, lines, and tokens [66]. The inputs to the iTrace Toolkit are: iTrace-Core data, iTrace Eclipse plugin data, and srcML archive files [67]. The fixation settings for generating fixations were: Fixation Filter: IVT [68], and the default settings for IVT: Velocity Threshold: 50, Duration (ms): 80. The iTrace Toolkit outputs a database containing fixation and gaze information. Then, we reviewed the screen recordings for each task and noted the x and y coordinates of each area of the screen (Figure 1, areas A-D). These coordinates varied for each task because we allowed each participant to configure their screen. We build on scripts from  [69], and the inputs are: databases from iTrace Toolkit, the “where” and “what” scores, screen region coordinates, and the self-reported confidence and difficulty scores.
3.7 Threats to Validity
@@
Our study has several threats to validity. One threat is that the selected bugs may not fully represent all C memory bugs. We mitigate this risk by choosing bugs as described in Section 3.3. A second threat is the varying detail in GitHub bug reports. To address this, we reviewed each report and made minor edits to ensure answering the “what” and “where” questions was non-trivial. Another threat is that all participants are graduate or undergraduate students, who may not fully represent industry professionals. To mitigate this, we selected only those with C experience and accepted only highly recommended undergraduates. Another threat is that some participants wore glasses and participants were not reminded to keep their heads still. However, the iTrace toolkit is designed to handle these concerns by using a conservative approach to identifying gazes and fixations.
@@
To address eye tracker drift [27] and occasional disconnections, we recalibrated the eye tracker between tasks and after any interruptions. An additional threat is that screen recordings often froze mid-session, affecting our ability to retrieve screen region coordinates and AI usage data. However, participants rarely changed window layouts after initial setup, and we used eye tracking data to confirm AI usage based on the amount of time the participant looked at the AI window. Also, the eye tracker was not tracking during Participant 11’s ladybug task, so this task is not included in eye tracking metrics, but the “where” and “what” scores for this task are included. Finally, Participant 11 did not complete the firefly task or answer the “what” and “where” questions within 30 minutes. Since we cannot infer their understanding, they received a score of 0, and we do not include this data when comparing successful versus unsuccessful tasks.
4 Experimental Results
4.1 RQ1: Task Outcomes
Overall, we found around a fifth of the tasks were completed successfully in terms of finding the correct bug location and identifying the correct reason for the bug. About 21% of the graded accuracy for “where” questions were 5, shown in Figure 2(f), which corresponds to a finding of “correct” (Table III). Also, only around 21% of the graded accuracy for “what” questions were 5, shown in Figure 2(c), which corresponds to a finding of “correctly” identifying the bug (Table II). However, these numbers rise to around 34% if we include answers clearly on the right track, only limited by time (i.e. accuracy scores of 4 or higher). These results show the difficulty of bug localization given low overall success, yet still point to the tractability of these bugs by people, considering that over half were at least on the right track despite only reading the bugs for a few minutes (57% of “where” accuracy scores were 3 or higher). Still, a minority did not show evidence that the problem would have been solved even given much more time — for only around 15% of tasks was there complete failure to find the correct location (score of 0 or 1 for “where” accuracy) and 23% for describing the problem (score of 0 or 1 for “what” accuracy), including total guesses or bug descriptions demonstrating key misunderstandings.
@@ 4.1 RQ1: Task Outcomes
Figure 2: Histograms of confidence, difficulty, and accuracy scores from all participants and tasks. Confidence and difficulty are self-reported by participants. Accuracy is graded on a rubric by the authors of this paper. “What” refers to the bug cause. “Where” refers to the bug location in code.
@@
What Confidence 1.000 -0.744 0.065 0.156 -0.208 0.099
Difficulty -0.744 1.000 0.083 -0.203 0.261 -0.038
Accuracy 0.065 0.083 1.000 0.414 -0.162 0.927
Where Confidence 0.156 -0.203 0.414 1.000 -0.616 0.458
Difficulty -0.208 0.261 -0.162 -0.616 1.000 -0.209
Accuracy 0.099 -0.038 0.927 0.458 -0.209 1.000
TABLE IV: Pearson correlation coefficient matrix. Significant values (except diagonal) at $\alpha=0.050$ are in bold.
@@
sway 6861 leak ladybug 1338 3.77 2.41 3.36 3.20 3.40 3.55
redis 8299 overflow stonefly 1424 3.14 3.41 2.73 3.23 2.95 3.14
sway 3577 double free hornet 1640 2.67 4.00 3.08 3.17 3.83 3.00
redis 9800 leak silverfish 900 3.50 3.50 1.50 3.50 3.00 1.50
redis 4527 overflow mantis 1284 3.00 4.00 2.20 2.00 4.25 2.20
sway 3300 double free spider 510 5.00 1.50 4.00 4.50 2.00 4.75
openssl 24892 leak firefly 1480 3.60 3.50 2.40 1.60 4.00 1.83
openssl 20278 double free weevil 930 1.75 4.00 3.88 3.75 3.25 3.50
TABLE V: Mean time required (in seconds) and mean scores per bug. Column “Issue Num” is the issue number in the project’s bug tracking system, included here for reproducibility. Column “Codename” is a name we used in our experiments and in this paper’s text for more memorable record-keeping.
Table IV shows Pearson correlations among all scores for all tasks and participants, with significant t-test results in bold. Notably the self-reported confidence and difficulty for “what” questions are not correlated to the rubric-graded accuracy for either question. In fact, there were around 13% of tasks for which participants self-reported high confidence (4 or 5) on “what” questions when graded accuracy scores were low (1 or 2). Still, confidence and difficulty scores are correlated, implying that participants are answering earnestly (random answers would tend to lead to uncorrelated responses). This observation may be due to the Dunning-Kruger effect, wherein people report high confidence in their knowledge despite weak performance [70].
Other correlations include: 1) “where” accuracy to “what” accuracy, which is not surprising considering that a correct rationale likely leads to a correct bug location, 2) “where” accuracy correlates to “where” confidence, implying that people are better judges of their performance of locating a bug, a finding supported by 3) “where” confidence is correlated with “what” accuracy, in that cases where people are confident in the location, they are more likely to give correct explanations (though, as noted above, not necessarily the other way around).
Overall participants reported being focused on the tasks at hand. In the attention question (prompted every five minutes), almost nine out of ten responses indicated attentiveness, as shown in Table VI. At a minimum, we interpret this finding to mean that we do not have evidence that participants were excessively distracted by the environment or other outside factors.
Option Percent
I was focused on the task. 87.68%
I was thinking about my performance on the task or how long it is taking. 9.85%
I was thinking about things unrelated to the task. 1.48%
I was distracted by sights or sounds in my environment. 0.99%
My mind was blank. 0.00%
TABLE VI: Self-Reported Attention Level Popup Options
TABLE VII: Gaze metrics for different bugs. Mean number of fixations for all bugs is 853 over 1188 seconds, or an average of 43 per minute.
Bug Fixation Count Regression Rate Fixation Duration (ms) Fixation Lines Fixation Functions
@@
ladybug 700 747 57 61 465 453 153 151 29 27
stonefly 1064 1149 51 57 449 437 199 199 32 33
hornet 1313 1560 52 58 412 426 248 215 30 12
mantis 687 798 43 39 484 413 201 209 18 22
firefly 1007 986 47 42 410 426 300 372 50 57
silverfish 638 638 42 42 402 402 200 200 29 29
spider 376 376 60 60 453 453 71 71 8 8
weevil 1037 636 60 64 422 419 163 131 13 12
TABLE VIII: Descriptive statistics of the number of top (most fixated on) functions that capture 75% of all fixations on functions
@@
ladybug 5.90 2 18 4.53 20% 11% 38% 8%
stonefly 3.91 2 7 1.51 16% 7% 43% 10%
hornet 10.17 2 46 17.57 31% 15% 43% 10%
mantis 3.60 1 7 2.30 21% 17% 33% 7%
firefly 7.17 1 12 4.75 14% 9% 22% 5%
silverfish 7.50 2 13 7.78 33% 25% 40% 11%
spider 1.50 1 2 0.71 25% 17% 33% 12%
weevil 3.25 3 4 0.50 34% 16% 60% 21%
all bugs 5.54 1 46 6.99 22% 7% 60% 12%
Figure 3: Bar charts showing the distribution of fixations over functions read by participants locating the bug firefly. Blue indicates the top functions that capture 75% of the total fixations on functions. Most reading effort is concentrated on a few functions.
4.2 RQ2: Visual Attention Patterns
We summarize broad patterns of visual attention in Tables VII - IX and Figures 3 - 5. Table VII shows: Most fixation counts are comparable to the average of 853, with the exception of bug spider which is much lower, at 376. The most likely explanation is that this bug is also the easiest, as Table V shows spider with the highest confidence, lowest difficulty, and highest accuracy. Unsurprisingly, the bug with the highest mean fixation count (hornet) was among the lowest confidence and tied for highest difficulty. In all cases, the mean and median fixation counts and regression rates per bug are relatively similar, indicating that outliers are not markedly affecting the mean. The mean fixation duration in milliseconds was stable throughout, with no clear association with bug difficulty.
TABLE IX: Descriptive statistics of the number of top (most fixated on) lines that capture 75% of all fixations on lines of code
@@ 4.2 RQ2: Visual Attention Patterns
ladybug 46.6 16 115 27.56 31% 19% 62% 13%
stonefly 49.82 19 109 26.37 29% 16% 68% 15%
hornet 96.50 52 207 60.41 37% 29% 62% 12%
mantis 71.80 33 123 36.51 37% 32% 44% 5%
firefly 107.50 11 164 65.05 33% 15% 40% 10%
silverfish 71.50 15 128 79.90 36% 36% 36% 1%
spider 22.50 10 35 17.68 30% 28% 33% 4%
weevil 42.50 29 60 15.46 29% 20% 35% 7%
all bugs 64.24 10 207 45.84 32% 15% 68% 11%
Figure 4: Bar charts showing the distribution of fixations over lines read by participants locating the bug firefly. Blue indicates the top lines that capture 75% of the total fixations on lines. Reading effort is distributed more widely than for functions.

### v2

@@ 3.6. Data Processing
Two authors (with 5 years and 14 years of teaching and grading experience respectively) graded the participants’ answers on a scale of 1-5 where 1 was very inaccurate and 5 was highly accurate. Attempts without answers received 0. A score of highly-accurate meant that the reported bug cause or location was identical to the one in the official issue or fixed by the official pull request or of equal plausibility as determined by the authors. For example, for ladybug, the location of the cause of the bug is line 48 (before fix): free(seat_device); because when seat_device is freed, access to seat_device->switch_device is lost, and the allocated memory for seat_device->switch_device is leaked. Notice that the bug cause location is not the fix location.
Table 2. Rubric for grading “what” questions.
@@
Table 3. Rubric for grading “where” questions.
@@
To minimize bias, we applied the rubrics shown in Tables 2 and 3 to decide scores. These rubrics reflect knowledge from related literature, the teaching experience of the authors, and the range of answers from participants. These rubrics are holistic, meaning that graders form an overall impression of the answer rather than assigning points for each criterion independently (Fitzgerald et al., 2013). The response characteristics that we looked for are similar to the requirements of the “isolation process” for debugging described in Hylton  et al. (Hylton et al., 2023), namely that the most complete answers have where the cause of the error is, the reasoning about which part of the code causes the error, and reasoning about how the error could have happened. Or, in other words, the most complete bug localizers know where, what, and why. We also reward coherent answers where the parts are “logically connected,” even if the answer is incorrect, over answers where the parts are not “logically connected” (Esche, 2024).
To resolve disagreements, we followed a discussion-based approach because it is recommended by related work for situations with a low tolerance for imperfect grading (Rieser and Lemon, 2011; Wood et al., 2018; Oortwijn et al., 2021). We followed the tested procedure from Oortwijn et al. (Oortwijn et al., 2021), which involves five steps. Step 1 is setting up the annotation task and guidelines. In our study, we use the rubrics in Tables 2 and 3, and each annotator annotates all participant responses. Step 2 is inter-rater conceptual alignment. During this step, we decide on the meaning of terms in our study like ”the cause of the bug” and ”the location of the bug.” These terms also had to be defined and explained to our participants, so we used the definitions and examples from the participant instructions. Step 3 is individual annotation when annotators score each response independently. Step 4 is annotation comparison. Step 5 is disagreement resolution. As noted in Oortwijn et al. (Oortwijn et al., 2021), annotators often find that disagreements are due to “task or guideline unclarity.” When this happens, annotators should clarify the guidelines. During this step, we discussed every disagreement and came to a consensus on what the final score should be, resulting in one final score for each response.
@@
To process and analyze the eye tracking data, we used the “Map Tokens” and “Generate Fixations” tools in iTrace Toolkit V0.2.2 to extract fixations, lines, and tokens (Behler et al., 2023). The inputs to the iTrace Toolkit are: iTrace-Core data, iTrace Eclipse plugin data, and srcML archive files (Collard et al., 2013). The fixation settings for generating fixations were: Fixation Filter: IVT (Andersson et al., 2017), and the default settings for IVT: Velocity Threshold: 50, Duration (ms): 80. The iTrace Toolkit outputs a database containing fixation and gaze information. Then, we reviewed the screen recordings for each task and noted the x and y coordinates of each area of the screen (Figure 2, areas A-D). These coordinates varied for each task because we allowed each participant to configure their screen. We build on scripts from Wallace  et al. (Wallace et al., 2025), and the inputs are: databases from iTrace Toolkit, the “where” and “what” scores, screen region coordinates, and the self-reported confidence and difficulty scores.
3.7. Threats to Validity
@@
Our study has several threats to validity. One threat is that the selected bugs may not fully represent all C memory bugs. We mitigate this risk by choosing bugs as described in Section 3.3. A second threat is the varying detail in GitHub bug reports. To address this, we reviewed each report and made minor edits to ensure answering the “what” and “where” questions was non-trivial. Another threat is that while most of the participants have professional experience, most participants are currently graduate or undergraduate students, who may not fully represent industry professionals. To mitigate this, we selected only those with C experience and accepted only highly recommended undergraduates. While our study was conducted with minimal distractions, the findings remain relevant because distraction effects are likely detectable and separable from bug-localization patterns. Another threat is that some participants wore glasses and participants were not reminded to keep their heads still. However, the iTrace toolkit is designed to handle these concerns by using a conservative approach to identifying gazes and fixations.
@@
To address eye tracker drift (Sharafi et al., 2020b) and occasional disconnections, we recalibrated the eye tracker between tasks and after any interruptions. An additional threat is that screen recordings often froze mid-session, affecting our ability to retrieve screen region coordinates and AI usage data. However, participants rarely changed window layouts after initial setup, and we used eye tracking data to confirm AI usage based on the amount of time the participant looked at the AI window. Also, the eye tracker was not tracking during Participant 11’s or Participant 18’s ladybug task, so these tasks are not included in eye tracking metrics. Finally, Participant 11 did not complete the firefly task and Participant 13 did not complete the mantis task. They did not answer the “what” and “where” questions within 30 minutes. Since we cannot infer their understanding, they received a score of 0, and we do not include this data when comparing successful versus unsuccessful tasks.
@@
Figure 3. Histograms of confidence, difficulty, and accuracy scores from all participants and tasks. Confidence and difficulty are self-reported by participants. Accuracy is graded on a rubric by the authors of this paper. “What” refers to the bug cause. “Where” refers to the bug location in code.:
4. Experimental Results
4.1. RQ1: Task Outcomes
Overall, we found around a fifth of the tasks were completed successfully in terms of finding the correct bug location and identifying the correct reason for the bug. About 18% of the graded accuracy for “where” questions were 5, shown in Figure 3(f), which corresponds to a finding of “correct” (Table 3). Also, only around 16% of the graded accuracy for “what” questions were 5, shown in Figure 3(c), which corresponds to a finding of “correctly” identifying the bug (Table 2). However, these numbers rise to around 38% if we include answers clearly on the right track, only limited by time (i.e. accuracy scores of 4 or higher). These results show the difficulty of bug localization given low overall success, yet still point to the tractability of these bugs by people, considering that about half were at least on the right track despite only reading the bugs for a few minutes (49% of “where” accuracy scores were 3 or higher). Still, a minority did not show evidence that the problem would have been solved even given much more time — for only around 16% of tasks was there complete failure to find the correct location (score of 0 or 1 for “where” accuracy) and 23% for describing the problem (score of 0 or 1 for “what” accuracy), including total guesses or bug descriptions demonstrating key misunderstandings.
Figure 4. Official fix for ladybug. See Figure 1 for the ladybug bug report. Expected bug cause location is line 48 (before fix): free(seat_device);.:
To illustrate the range of participant answers, consider the bug ladybug, a memory leak of a variable called seat_device->switch_device. See Figures 1 and 4 for more information about ladybug. An example of an answer that received a “what” score of 1:
“The bug is that the calloc function is making a recursive call to the function sway_switch_create and is sending the wrong parameters. We are allocating possibly more memory than needed. The parameters are wrong, need to redefine before the call and do not send it the same function.”
This answer received a “what” score of 1 because it is not related to the memory leak bug. An example of an answer that received a “what” score of 3:
“The cause of the bug is a calloc call that is not properly freed. The memory allocation that should be freed occurs on line 9.”
This answer received a score of 3 because it is incomplete. While the participant does identify which memory is leaked, the participant does not identify where or why. An example of an answer that received a “what” score of 5:
“In file seat.c, line no 881, seat_configure_switch function to configure switch. In line 815, it creates a switch, if it does not exists. To do so, it calls sway_switch_create method from switch.c file. In this file, at line no 9, it allocates memory for the switch. However, when seat_device_destroy method is called to destroy the seat_device in seat.c file, in this method from line 37 to 49, the allocated switch is not destroyed by calling sway_switch_destroy method from switch.c file. It can cause memory leak issue.”
This answer received a score of 5 because it identifies the cause of the bug by explaining why the memory leak occurs.
Table 4. Pearson correlation coefficient matrix. Significant values (except diagonal) at $\alpha=0.050$ are in bold.
@@ 4.1. RQ1: Task Outcomes
What
Confidence 1.000 -0.740 0.291 0.542 -0.452 0.272
Difficulty -0.740 1.000 -0.183 -0.452 0.556 -0.174
Accuracy 0.291 -0.183 1.000 0.376 -0.339 0.930
Where
Confidence 0.542 -0.452 0.376 1.000 -0.643 0.324
Difficulty -0.452 0.556 -0.339 -0.643 1.000 -0.280
Accuracy 0.272 -0.174 0.930 0.324 -0.280 1.000
Table 4 shows Pearson correlations among all scores for all tasks and participants, with significant t-test results in bold. Notably the self-reported difficulty for “what” questions is not correlated to the rubric-graded accuracy for either question. Still, confidence and difficulty scores are correlated, implying that participants are answering earnestly (random answers would tend to lead to uncorrelated responses). Other correlations include: 1) “where” accuracy to “what” accuracy, which is not surprising considering that a correct rationale likely leads to a correct bug location, 2) “where” accuracy correlates to “where” confidence and “where” difficulty, implying that people are better judges of their performance of locating a bug, a finding supported by 3) “where” confidence is correlated with “what” accuracy, in that cases where people are confident in the location, they are more likely to give correct explanations.
Overall participants reported being focused on the tasks at hand. In the attention question (prompted every five minutes), nine out of ten responses indicated attentiveness, as shown in Table 5. While we cannot know peoples’ mental states with 100% certainty, one possible explanation for the high level of reported attentiveness is the duration of the tasks (30 minutes per bug). Studies with longer tasks are more likely to lead to fatigue and inattentiveness. At a minimum, we interpret this finding to mean that we do not have evidence that participants were excessively distracted by the environment or other outside factors. We do not analyze the eye tracking data in relation to the attentiveness survey due to the scarcity of data points showing inattentiveness.
Table 5. Self-Reported Attention Level Popup Options
Option Percent
I was focused on the task. 91.03%
I was thinking about my performance on the task or how long it is taking. 6.06%
I was thinking about things unrelated to the task. 1.58%
I was distracted by sights or sounds in my environment. 0.79%
My mind was blank. 0.00%
Table 6. Mean time required (in seconds) and mean scores per bug. Column “Issue Num” is the issue number in the project’s bug tracking system, included here for reproducibility. Column “Codename” is a name we used in our experiments and in this paper’s text for more memorable record-keeping.
@@
sway 6861 leak ladybug 1330 3.80 2.48 3.20 3.50 2.95 3.20
redis 8299 overflow stonefly 1426 3.02 3.31 2.67 3.17 3.38 2.67
sway 3577 double free hornet 1536 2.88 4.00 3.00 3.13 4.00 3.13
redis 4527 overflow mantis 1159 2.89 3.61 1.78 2.31 3.94 2.22
openssl 24892 leak firefly 1259 3.00 3.79 1.86 1.71 4.14 1.71
redis 9800 leak silverfish 1241 2.14 4.14 1.71 2.29 3.86 1.86
sway 3300 double free spider 645 4.67 1.61 4.67 4.28 1.83 4.78
openssl 20278 double free weevil 933 2.39 3.67 3.67 3.11 3.28 3.44
Table 7. Gaze metrics on code for different bugs. Mean number of fixations on code for all bugs is 838 over 1191 seconds, or an average of 42 per minute.
Bug Fixation Count Regression Rate Duration (ms) Lines Functions
@@
ladybug 773 756 62 62 450 440 153 146 27 23
stonefly 1024 1022 52 56 447 450 198 172 31 30
hornet 1250 1560 51 54 412 426 239 215 25 12
mantis 658 695 44 45 474 433 182 180 17 18
firefly 890 937 44 42 403 416 260 302 43 43
silverfish 588 591 43 47 380 393 170 135 21 12
spider 460 425 55 56 423 416 103 102 11 12
weevil 1062 770 67 67 435 422 125 95 8 6
(a) Bar charts showing the distribution of fixations over functions read by participants locating the bug firefly. Blue indicates the top functions that capture 75% of the total fixations on functions. Reading effort is concentrated on a few functions.:
(b) Bar charts showing the distribution of fixations over lines read by participants locating the bug firefly. Blue indicates the top lines that capture 75% of the total fixations on lines. Reading effort is distributed more widely than for functions.:
Figure 5. Bar charts showing the distribution of fixations over functions and lines of code.:
Table 8. Descriptive statistics of the number of most fixated on functions that capture 75% of all fixations on functions
@@
ladybug 5.4 1 18 4.3 20% 8% 38% 7%
stonefly 4.2 2 11 2.2 18% 7% 67% 14%
hornet 8.5 2 46 15.2 34% 15% 60% 14%
mantis 3.3 1 7 1.8 20% 11% 33% 6%
firefly 6.4 1 12 4.3 17% 9% 33% 8%
silverfish 4.7 2 13 4.3 31% 10% 67% 18%
spider 2.2 1 4 0.8 24% 15% 40% 9%
weevil 2.8 1 4 0.8 51% 16% 100% 26%
all bugs 4.6 1 46 5.4 25% 7% 100% 17%
Table 9. Descriptive statistics of the number of most fixated on lines that capture 75% of all fixations on lines of code
@@
ladybug 43.6 14 141 34.7 28% 17% 62% 10%
stonefly 51.6 12 129 29.3 28% 16% 68% 12%
hornet 91.9 44 207 54.9 38% 29% 62% 11%
mantis 61.5 19 123 29.5 34% 22% 44% 7%
firefly 94.3 11 164 63.6 35% 15% 47% 10%
silverfish 63.1 11 168 61.0 35% 10% 53% 14%
spider 33.6 10 61 17.8 32% 25% 36% 4%
weevil 36.3 23 60 11.8 31% 20% 42% 6%
all bugs 56.0 10 207 41.7 31% 10% 68% 10%
4.2. RQ2: Visual Attention Patterns
We summarize broad patterns of visual attention in Tables 7 - 9 and Figures 5(a) - 6. Table 7 shows a considerable spread in fixation count among the bugs. The bug spider has the lowest fixation count, at 460. The most likely explanation is that this bug is also the easiest, as Table 6 shows spider with the highest confidence, lowest difficulty, and highest accuracy. Unsurprisingly, the bug with the highest mean fixation count (hornet) was among the highest difficulty. However, fixation count is not a perfect indicator of difficulty. For example, silverfish has a relatively low fixation count, but a high difficulty. This may be explained when looking at the average time for silverfish. This high time but low fixation count on code indicates that participants spent a lot of time reading the bug report. The silverfish bug report was not particularly long, but participants may have had difficulty interpreting the report. In all cases, the mean and median fixation counts and regression rates per bug are relatively similar, indicating that outliers are not markedly affecting the mean. In this work, “regression rate” is the percentage of fixations where the token fixated on was fixated on more than once (Wallace et al., 2025). The mean fixation duration in milliseconds was stable throughout, with no clear association with bug difficulty.
A novel observation of this paper is that people read several functions outside of the one where the bug is located (a mean of 22), though visual attention is concentrated among relatively few.
Key Finding 1: Table 8 shows that programmers spend 75% of their visual effort on 25% of the functions that they view.